# Follow the Flow: On Information Flow Across Textual Tokens in Text-to-Image Models

### v1

@@ A.5.2 SDXL-Turbo
Our analysis reveals that SDXL-Turbo, which uses the CLIP text encoder, behaves markedly differently from FLUX, which relies on the encoder in the encoder-decoder T5-XXL. In SDXL-Turbo, the text encoder is a causal language model, meaning each token’s encoding is influenced only by its preceding tokens during the encoding process. We repeated our in-item information flow experiments using SDXL-Turbo. Our first observation is that most generated images are either abstract or unrelated to the intended lexical items. According to our analysis, 55% of lexical items in SDXL-Turbo lack any representative token (compared to just 11% in FLUX). Moreover, when a representative token is present in CLIP, it is typically the final token of the lexical item (see 10). This is aligned with the unidirectional encoding of the model. Another phenomenon we observe—consistent with CLIP’s training objective—is the unusually dominant role of the end-of-sequence (EOS) token. Images generated from the EOS token often encapsulate nearly the full semantic content of the prompt. In our evaluation, 62% of EOS-generated images matched the prompt ( compared to 73% when using the full prompt). We believe this also causes our intervention method to be less effective, since when we interpret a single token, we patch all other tokens, including the EOT token—which usually contains a lot of information—with tokens derived from an empty prompt (see 10).

### v2

@@ A.3.2 SDXL-Turbo
Our analysis reveals that SDXL-Turbo, which uses the CLIP text encoder, behaves markedly differently from FLUX, which relies on the encoder in the encoder-decoder T5-XXL. In SDXL-Turbo, the text encoder is a causal language model, meaning each token’s encoding is influenced only by its preceding tokens during the encoding process.
Figure 10: Token-level image generation using Sana on the prompt “a black baseball hat with a flame decal on it”. Representative information often resides in the last token of each item (e.g., “hat”, “decal”), consistent with Sana’s causal encoding. Contextual influence is one-directional, with earlier tokens shaping later ones.
We repeated our in-item representation experiments using SDXL-Turbo. Our first observation is that most generated images are either abstract or unrelated to the intended lexical items. According to our analysis, 55% of lexical items in SDXL-Turbo lack any representative token (compared to just 11% in FLUX). Moreover, when a representative token is present in CLIP, it is typically the final token of the lexical item (see Fig.˜9). This is aligned with the unidirectional encoding of the model.
@@
Another phenomenon we observe—consistent with CLIP’s training objective—is the unusually dominant role of the end-of-sequence (EOS) token. Images generated from the EOS token often encapsulate nearly the full semantic content of the prompt. In our evaluation, 62% of EOS-generated images matched the prompt ( compared to 73% when using the full prompt). We believe this also causes our intervention method to be less effective, since when we interpret a single token, we patch all other tokens, including the EOT token—which usually contains a lot of information—with tokens derived from an empty prompt (see Fig.˜9).
A.3.3 Sana (Gemma-based encoder)
We also evaluate our methodology on the Sana model (Xie et al., 2024b), which uses a Gemma-based autoregressive language model as the text encoder. These results help validate the generality of our findings across architectures with differing encoding strategies.
For in-item information flow, we observe that Sana produces fewer multi-token lexical items due to its larger vocabulary relative to T5 and CLIP. In cases where items do consist of multiple tokens, we find that only the last token typically acts as a representative token—a behavior aligned with the unidirectional nature of autoregressive encoders.
In our cross-item information flow analysis, we find that 17.45% of item pairs in Sana exhibit contextual information flow, compared to 10.45% in FLUX. This increase likely stems from the one-sided (forward-only) nature of Sana’s encoding. For example, in the prompt “a black baseball hat with a flame decal on it”, we find contextual influence such as “black baseball” affecting “hat” and “flame” affecting “decal”, but not the reverse. See Fig.˜10 for illustrations.
Overall, while Sana shows a slightly higher rate of contextual influence than FLUX, it preserves many of the key structural insights found in our primary analysis—particularly the sparsity and location of representative tokens. Unlike CLIP, which encodes substantial information in the EOS token, Sana lacks such artifacts, suggesting that EOS-related effects are not fundamental to autoregressive models more generally.
A.4 Inter-Item Information Flow Framework
To assess whether one lexical item encodes information about others in the same prompt, we conduct the following experiment.
Given a prompt, we isolate each lexical item one at a time. First, we encode the full prompt using the text encoder. Then, for a given lexical item, we apply our patching method (Section˜2.1) to mask the representations of all other tokens—leaving only the contextualized representation of the target item intact. Formally, for a lexical item with token indices $S\subset\{1,\ldots,N\}$, we construct a patched sequence $\tilde{t}_{1},\ldots,\tilde{t}_{N}$ in which only the tokens in $S$ retain their original contextualized representations. We then generate an image from this modified sequence, capturing what information is encoded in the selected item’s contextualized form.
We repeat this process for the same item in an uncontextualized setting: we encode and generate an image using only that lexical item in isolation, without the rest of the prompt. This allows us to distinguish between information inherently present in the item’s encoding and information introduced by context.
To measure influence between items, we use Qwen2-VL to check whether a second item $y$ appears in the image generated from a first item $x$, both in the contextualized and uncontextualized versions. Influence is considered True if $y$ appears in the image generated from contextualized $x$, but not from uncontextualized $x$. This comparison ensures that the observed presence of $y$ is attributable to contextual information flow during encoding, rather than coincidental co-occurrence or inherent correlation.
This setup enables us to detect breaches of lexical segmentation—i.e., cases where one lexical item encodes visual information belonging to another—quantifying interdependence across items within the prompt.
See Fig.˜11 for an illustration of the procedure.
A.5 Predicting Redundant Tokens Without Image Generation
Identifying which tokens are representative can improve image generation, as demonstrated in Fig.˜4 (right). Moreover, many existing T2I methods operate at the token level, typically treating all of a lexical item’s tokens equally (Chefer et al., 2023a; Dahary et al., 2024). However, our findings suggest that only a subset of these tokens contribute meaningfully to the representation.
Yet, identifying representative tokens via image generation is computationally expensive. To address this, we explore alternatives for estimating token importance without image-level interventions.
First, we compute the character-level edit distance between a lexical item (e.g., “giraffe”) and its constituent tokens (e.g., “gir”, “a”, “ffe”). We observe a negative correlation between edit distance and token representativeness (Pearson correlation of 0.44-0.44), suggesting that tokens more orthographically distant from the original item are less likely to be representative. While informative, this signal alone is insufficient for accurate prediction.
Instead, we train a $k$-Nearest Neighbors (k=5k{=}5) classifier using Euclidean distance over the token representations to predict redundancy. Using 6,966 token instances (spanning 4,864 unique lexical items), each labeled for whether the token is representative according to a VLM, our classifier achieves 92% precision in identifying non-representative tokens. This result highlights the feasibility of lightweight approaches for filtering redundant tokens without resorting to costly generation procedures. See Section˜A.1.2 for classifier details.